# Can the neutrinos from TXS 0506+056 have a coronal origin?

### v1

@@ 1 Introduction
These recent developments have demonstrated that non-jetted AGN may be efficient emitters of neutrinos. This naturally raises the question of the respective contributions of the corona and the jet to the neutrino signal expected from blazars. TXS 0506+056, undoubtedly, stands out as an ideal source for addressing this question. Recently, Khatee Zathul et al. (2024); Yang et al. (2024) proposed that multi-messenger observations of TXS 0506+056 might be interpreted by neutrino production in the core region of the AGN, namely a compact region close to the black hole, such as the corona or the inner accretion flow, rather than the blazar jet as previously assumed. This would have consequences not only on the specific question of the TXS 0506+056 neutrinos, but more broadly on the blazar contribution to the diffuse neutrino sky compared to non-jetted AGN.
In this Letter, we re-examine the corona hypothesis for neutrino production in TXS 0506+056. We use observational properties of the blazar to infer its accretion-driven X-ray luminosity. Within the context of the scenario of Fiorillo et al. (2024b) (hereby denoted F24), in which non-thermal particles are energized within a magnetic reconnection layer close to the central black hole, we construct a model for coronal neutrino production consistent with the energetics and the X-ray luminosity in the corona. By comparing to the potential emission from a jet (blazar), we then critically assess the possibility that the observed neutrinos from TXS 0506+056 come from the corona.
@@ 2 Coronal model of neutrino emission
We summarize the main properties of the coronal region and the mechanism by which non-thermal particles may be accelerated and interact within it. The corona is assumed to be a compact region located close to the central supermassive black hole, with dimensions up to a few tens of gravitational radii $r_{g}=GM/c^{2}$, where $G$ is gravitational constant, $c$ is the speed of light, and $M$ is the mass of the central supermassive black hole. For TXS 0506+056, we use $M\simeq 3\times 10^{8}\;M_{\odot}$ (Padovani et al., 2019), so $r_{g}\simeq 4.4\times 10^{13}$ cm. The corona itself is permeated by a dense X-ray field, generated by Comptonization of low-energy photons – either from the accretion disk or internal to the corona – from energetic leptons (Sunyaev & Truemper, 1979; Sunyaev & Titarchuk, 1980; Haardt & Maraschi, 1991, 1993) or from the bulk motion of turbulent plasma (Groselj et al., 2023) or bulk motion of plasmoids in reconnection layers (Beloborodov, 2017; Sironi & Beloborodov, 2020; Sridhar et al., 2021, 2023). The X-ray field also plays a key role in non-thermal dynamics, since it acts as a target for $p\gamma$ and $\gamma\gamma$ interactions, rendering the corona opaque to high-energy gamma rays and an efficient converter of proton energy into neutrinos and electromagnetic radiation. The optical thickness of the corona to said processes depends crucially on the size of the corona and especially on the mechanism by which non-thermal particles are accelerated.
@@
We first discuss the properties of the proton injection spectrum. PIC simulations [^1] These do not include cooling processes for relativistic protons. show that the differential proton number density can be described by a broken power law, $dn_{p}/dE_{p}\propto E_{p}^{-1}$ for $E\leq E_{p,\rm br}\simeq\sigma_{p}m_{p}c^{2}$ and $dn_{p}/dE_{p}\propto E_{p}^{-s}$ for $E>E_{p,\rm br}$. Protons accelerated by the reconnection process are in rough energy equipartition with the reconnecting magnetic field. The post-break power-law index $s$ ranges from $2$ to $3$ depending on the strength of the non-reconnecting component of the magnetic field, known as the guide field. A strong guide field ($\gtrsim 0.5B$) renders acceleration less efficient and therefore produces a soft power law with $s\simeq 3$, while $s\simeq 2$ for weak guide fields. At an energy large enough that the acceleration rate is comparable with radiative cooling rate, which we define as $E_{p,\rm rad}$, the spectrum exhibits a cutoff. This summarizes the properties of the proton injection spectrum in the downstream region.
@@ 4 Multi-messenger emission from TXS 0506+056
Here we normalize the radius $R$ in units of $R_{13.8}=R/1.5r_{g}$. The magnetic energy density is then related to the X-ray photon energy density as $u_{B}=u_{X}/(6\beta_{\rm rec}\eta_{X})\approx 3u_{X}\beta^{-1}_{\rm rec,-1}\eta^{-1}_{X,-0.3}$.
@@
In all three cases, for the proton injection spectrum we consider a high-guide-field (HGF) case, in which $s=3$, and we choose $\sigma_{p}=10^{7}$ ($E_{p,\rm br}\simeq 9.4$ PeV). This ensures that neutrinos can be produced up to PeV energies, as observed by IceCube. Neutrino production at PeV can also be obtained by assuming a low-guide-field (LGF) scenario, in which $s=2$, and we choose $\sigma_{p}=10^{5}$ and ; results for this alternative case are very similar to the HGF one and we show them in Appendix B.
@@ 5 Discussion
The discrepancy with previous works (Khatee Zathul et al., 2024; Yang et al., 2024) primarily comes from the stringent observational constraints on the coronal X-ray luminosity. While Khatee Zathul et al. (2024) explicitly addresses the corona as neutrino production site, Yang et al. (2024) refers instead to the accretion flow in the inner regions close to the supermassive black hole. However, the qualitative features of the model are essentially very similar. Yang et al. (2024) assumes an acceleration timescale comparable with the gyroradius of the proton (divided by $c$); while this is not associated with a specific acceleration mechanism, such a rapid energization cannot happen via stochastic acceleration in turbulence, and must instead be powered by magnetic reconnection. Therefore, our generic constraints on the X-ray and the proton luminosity apply. In particular, the proton luminosity assumed in Yang et al. (2024), of the order of $L_{p}\sim 10^{46}$ erg/s, is about two orders of magnitude higher than $L_{p}$ in our HL scenario, and is therefore difficult to reconcile with the expected near-equipartition between X-rays and protons within a reconnection layer.
@@ Appendix B Low-guide-field scenario
In this section, we show for completeness how the results in the main text when an injection scenario with low guide field is considered. In this case, protons are injected with a $\sigma_{p}=10^{5}$, corresponding to a break energy of about $100$ TeV, and have a power-law index $s=2$ at higher energies.

### v2

@@ 1 Introduction
These recent developments have demonstrated that non-jetted AGN may be efficient emitters of neutrinos. This naturally raises the question of the respective contributions of the corona and the jet to the neutrino signal expected from blazars. TXS 0506+056, undoubtedly, stands out as an ideal source for addressing this question. Kimura & Toma (2020) concluded that the neutrino luminosity from a hot accretion flow in TXS 0506+056 should be smaller than about $10^{44}$ erg/s for a magnetically arrested disk (MAD) with sub-Eddington accretion. However, recently, Khatee Zathul et al. (2024); Yang et al. (2024) proposed that multi-messenger observations of TXS 0506+056 might be interpreted by neutrino production in the core region of the AGN, namely a compact region close to the black hole, such as the corona or the inner accretion flow, rather than the blazar jet as previously assumed. Yang et al. (2024), using a modeling similar to Kimura & Toma (2020), concludes that the 10-years-averaged neutrino signal from TXS 0506+056might indeed be produced in the hot accretion flow, and that even the 2014/15 neutrino flare might be explained if temporary super-Eddington accretion is assumed. A core neutrino production would have consequences not only on the specific question of the TXS 0506+056 neutrinos, but more broadly on the blazar contribution to the diffuse neutrino sky compared to non-jetted AGN.
In this work, we re-examine the corona hypothesis for neutrino production in TXS 0506+056. We use observational properties of the blazar to infer its accretion-driven X-ray luminosity. Within the context of the scenario of Fiorillo et al. (2024b) (hereby denoted F24), in which non-thermal particles are energized within a magnetic reconnection layer close to the central black hole, we construct a model for coronal neutrino production consistent with the energetics and the X-ray luminosity in the corona. By comparing to the potential emission from a jet (blazar), we then critically assess the possibility that the observed neutrinos from TXS 0506+056 come from the corona.
@@ 2 Coronal model of neutrino emission
We summarize the main properties of the coronal region and the mechanism by which non-thermal particles may be accelerated and interact within it. The corona is assumed to be a compact region located close to the central supermassive black hole, with dimensions up to a few tens of gravitational radii $r_{g}=GM/c^{2}$, where $G$ is gravitational constant, $c$ is the speed of light, and $M$ is the mass of the central supermassive black hole. For TXS 0506+056, we use $M\simeq 3\times 10^{8}\;M_{\odot}$ (Padovani et al., 2019), so $r_{g}\simeq 4.4\times 10^{13}$ cm. The corona itself is permeated by a dense X-ray field, generated by Comptonization of low-energy photons – either from the accretion disk or internal to the corona – from energetic leptons (Sunyaev & Truemper, 1979; Sunyaev & Titarchuk, 1980; Haardt & Maraschi, 1991, 1993) or from the bulk motion of turbulent plasma (Grošelj et al., 2024) or bulk motion of plasmoids in reconnection layers (Beloborodov, 2017; Sironi & Beloborodov, 2020; Sridhar et al., 2021, 2023). The X-ray field also plays a key role in non-thermal dynamics, since it acts as a target for $p\gamma$ and $\gamma\gamma$ interactions, rendering the corona opaque to high-energy gamma rays and an efficient converter of proton energy into neutrinos and electromagnetic radiation. The optical thickness of the corona to said processes depends crucially on the size of the corona and especially on the mechanism by which non-thermal particles are accelerated.
@@
We first discuss the properties of the proton injection spectrum. PIC simulations [^1] These do not include cooling processes for relativistic protons. show that the differential proton number density can be described by a broken power law, $dn_{p}/dE_{p}\propto E_{p}^{-1}$ for $E\leq E_{p,\rm br}\simeq\sigma_{p}m_{p}c^{2}$ and $dn_{p}/dE_{p}\propto E_{p}^{-s}$ for $E>E_{p,\rm br}$. Protons accelerated by the reconnection process are in rough energy equipartition with the reconnecting magnetic field. The post-break power-law index $s$ ranges from $2$ to $3$ depending on the strength of the non-reconnecting component of the magnetic field, known as the guide field. A strong guide field ($\gtrsim 0.5B$) renders acceleration less efficient and therefore produces a soft power law with $s\simeq 3$, while $s\simeq 2$ for weak guide fields. A strong guide field is favored from the IceCube detection of NGC 1068 neutrinos (Fiorillo et al., 2024b) and has been observed in the reconnection layers formed at the black hole jet sheath Sridhar et al. (2025); for our purposes, however, the difference between the two cases is not essential, as we will discuss. At an energy large enough that the acceleration rate is comparable with radiative cooling rate, which we define as $E_{p,\rm rad}$, the spectrum exhibits a cutoff. This summarizes the properties of the proton injection spectrum in the downstream region.
@@ 4 Multi-messenger emission from TXS 0506+056
Here we normalize the radius $R$ in units of $R_{13.8}=R/1.5r_{g}$. The magnetic energy density is then related to the X-ray photon energy density as $u_{B}=u_{X}/(6\beta_{\rm rec}\eta_{X})\approx 3u_{X}\beta^{-1}_{\rm rec,-1}\eta^{-1}_{X,-0.3}$. The corresponding magnetization parameter for the lepton pairs $\sigma_{e}=B^{2}/(4\pi n_{e}m_{e}c^{2})$ is in the range $10-100$ for the inferred values of $L_{X}$ and $n_{e}$ (assuming a corona with $\tau_{T}=0.5$).
@@
In all three cases, for the proton injection spectrum we consider a high-guide-field (HGF) case, in which $s=3$, and we choose $\sigma_{p}=10^{7}$ ($E_{p,\rm br}\simeq 9.4$ PeV). This ensures that neutrinos can be produced up to PeV energies, as observed by IceCube. Neutrino production at PeV can also be obtained by assuming a low-guide-field (LGF) scenario, in which $s=2$, and we choose $\sigma_{p}=10^{5}$; results for this alternative case are very similar to the HGF one and we show them in Appendix B. These two cases, which are consistent with a pair-dominated corona ($\sigma_{p}\gg\sigma_{e}$), are meant to show the impact of $\sigma_{p}$, one of the main unconstrained parameters of our model; see Karavola et al. (2024) for a systematic discussion on the effects of $\sigma_{p}$ on the neutrino spectra.
@@ 5 Discussion
The discrepancy with previous works (Khatee Zathul et al., 2024; Yang et al., 2024) primarily comes from the stringent observational constraints on the coronal X-ray luminosity. While Khatee Zathul et al. (2024) explicitly addresses the corona as neutrino production site, Yang et al. (2024) refers instead to the accretion flow in the inner regions close to the supermassive black hole. However, the qualitative features of the model are essentially very similar. Yang et al. (2024) assumes an acceleration timescale comparable with the gyroradius of the proton (divided by $c$); while this is not associated with a specific acceleration mechanism, such a rapid energization cannot happen via stochastic acceleration in turbulence, and must instead be powered by magnetic reconnection. Therefore, our generic constraints on the X-ray and the proton luminosity apply. In particular, the proton luminosity assumed in Yang et al. (2024), of the order of $L_{p}\sim 10^{46}$ erg/s, is about two orders of magnitude higher than $L_{p}$ in our HL scenario, and is therefore difficult to reconcile with the expected near-equipartition between X-rays and protons within a reconnection layer. In addition, if the accretion flow is responsible for neutrino production, then super-Eddington accretion may raise additional tension with acceleration of protons at large energies, since the increased density may turn the plasma collisional and hinder efficient acceleration. However, it is difficult to substantiate these statements without a concrete estimate of the acceleration rate, which depends on the assumed mechanism. Our arguments, based on energetics, are independent of these choices.
@@ Appendix B Low-guide-field scenario
In this section, we show for completeness in Fig. 4 how the results in the main text when an injection scenario with low guide field is considered. In this case, protons are injected with a $\sigma_{p}=10^{5}$, corresponding to a break energy of about $100$ TeV, and have a power-law index $s=2$ at higher energies.